A and B are indirectly proportional to each other. When B = 120, A = 30. What is the value of A when B = 160.
Found 2 solutions by mukhopadhyay, sarah_adam:
Answer by mukhopadhyay(490)

The question should read as "A and B are inversely proportional to each other"
A=k/B (because they are inversely proportional to each other)
30=k/120
=> k = 30*120
A = 30*120/B
When B = 160, A = 30*120/160 = 360/16 = 45/2;

Answer by sarah_adam(201)

When A and B are inversely proportional A = K/B where K = constant
Now when A = 30 and B = 120
then K = A*B
K = 30*120 = 3600
No when B = 160
A = 3600/160 = 22.5